OpenStudy (anonymous):

Simplify the following expression: 2x - 4y + 3x2 + 7y - 16x

OpenStudy (anonymous):

Combine the like terms: 2x - 16x = -14x and -4y + 7y = 3y 3x^2 - 14x + 3y

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

2x - 4y + 3x^2 + 7y - 16x = 3x^2 - 14x + 3y.

OpenStudy (anonymous):

ahhhh okay

OpenStudy (anonymous):

ty